OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4. 1). 1 2).-1 3).-5 4).-25

OpenStudy (fibonaccichick666):

so any ideas on how to begin? What have you tried so far?

OpenStudy (anonymous):

I came up with -5. But i have a feeling its completely wrong.

OpenStudy (fibonaccichick666):

ok, so do you want to learn how to check that with me?

OpenStudy (fibonaccichick666):

its not hard

OpenStudy (fibonaccichick666):

:)

OpenStudy (anonymous):

3(1 + x) = -2(x + -1) + -4 (1 * 3 + x * 3) = -2(x + -1) + -4 (3 + 3x) = -2(x + -1) + -4

OpenStudy (anonymous):

3 + 3x = -2(-1 + x) + -4 3 + 3x = (-1 * -2 + x * -2) + -4 3 + 3x = (2 + -2x) + -4

OpenStudy (fibonaccichick666):

that is correct so far

OpenStudy (fibonaccichick666):

good to there too

OpenStudy (fibonaccichick666):

I like that you use plus a negative 1

OpenStudy (anonymous):

I got the answer I just made a little mistake. Thank you for making me realize it. The answer is -1.

OpenStudy (fibonaccichick666):

no problem, and if you ever want a quick check on your own just substitute what you get back in for your variable, then you know if you made a math error